# SegQuant: A Semantics-Aware and Generalizable Quantization Framework for Diffusion Models

### v2

@@ Introduction
To quantify this asymmetry, we analyzed activation outputs channel-wise across several representative layers from SD3.5-ControlNet using the COCO dataset (averaged over 30 timesteps). As shown in Table 1, a large proportion of channels consistently exhibit negative values. For instance, over 60%-70% of channels in certain AdaNorm and FFN layers remain predominantly negative during inference. However, conventional PTQ methods typically apply a single global scale (either symmetric or asymmetric) across the entire activation range. Due to the large spread of positive values, compression of the narrower negative range ($[-0.3,0]$) is often too aggressive. Even with asymmetric quantization, which shifts the zero-point, the uniform distribution of bins leads to poor resolution for negative values. Figure 7 shows that modifying only the negative activations causes subtle but perceptible detail loss, particularly in texture and background consistency. This indicates that negative activations are not residual noise but semantically meaningful components. Although overall quality remains largely intact, this underscores preserving small negative values for high-fidelity generation.
@@
SD3.5-DiT 8/8(int) Q-Diffusion $169.59$ $-2.072$ $0.690$ $7.92$ $0.295$
PTQD $26.53$ $0.309$ $0.520$ $10.20$ $0.417$
PTQ4DiT $16.46$ $0.752$ $0.426$ $12.18$ $0.532$
TAC $17.78$ $0.702$ $0.440$ $11.99$ $0.520$
SegQuant-A $13.90$ $0.924$ $0.384$ $12.78$ $0.563$
SegQuant-G $12.37$ $0.859$ $0.383$ $12.83$ $0.564$
4/8(int) PTQ4DiT $62.98$ $-0.190$ $0.577$ $10.06$ $0.429$
SVDQuant $20.58$ $0.725$ $0.456$ $11.76$ $0.523$
SegQuant-G $20.22$ $0.762$ $0.453$ $11.69$ $0.521$
SDXL-UNet 8/8(int) Q-Diffusion $72.81$ $-1.618$ $0.567$ $11.97$ $0.360$
PTQ4DiT $7.70$ $0.736$ $0.191$ $19.66$ $0.691$
SegQuant-A $6.39$ $0.775$ $0.141$ $21.33$ $0.742$
SegQuant-G $6.19$ $0.764$ $0.134$ $21.60$ $0.750$
8/8(fp) Q-Diffusion $5.14$ $0.897$ $0.093$ $24.31$ $0.827$
SegQuant-A $5.12$ $0.901$ $0.093$ $24.28$ $0.827$
SegQuant-G $4.83$ $0.903$ $0.082$ $24.84$ $0.838$
FLUX-DiT 8/8(int) Q-Diffusion $9.41$ $0.732$ $0.299$ $15.87$ $0.633$
PTQ4DiT $11.91$ $0.630$ $0.325$ $15.36$ $0.611$
SegQuant-A $5.89$ $0.835$ $0.150$ $19.84$ $0.770$
SegQuant-G $5.56$ $0.822$ $0.138$ $20.32$ $0.782$
4/8(int) PTQ4DiT $30.09$ $-0.039$ $0.540$ $12.09$ $0.540$
SVDQuant $7.94$ $0.783$ $0.232$ $17.29$ $0.697$
SegQuant-G $7.78$ $0.789$ $0.225$ $17.48$ $0.702$
@@ Appendix B More Experimental Details and Results
SD3.5-DiT 8/8(int) Q-Diffusion $218.04$ $-2.197$ $0.395$ $16.02$ $0.704$ $8.21$ $0.388$
PTQD $15.62$ $0.822$ $0.434$ $16.47$ $0.454$ $10.79$ $0.523$
PTQ4DiT $14.21$ $0.912$ $0.458$ $16.37$ $0.403$ $12.69$ $0.593$
TAC $16.89$ $0.863$ $0.458$ $16.31$ $0.417$ $12.53$ $0.583$
SegQuant-A $10.90$ $1.020$ $0.467$ $16.35$ $0.362$ $13.27$ $0.618$
SegQuant-G $11.06$ $0.991$ $0.457$ $16.36$ $0.376$ $13.14$ $0.601$
4/8(int) PTQ4DiT $92.36$ $0.085$ $0.411$ $16.51$ $0.570$ $10.83$ $0.504$
SVDQuant $25.48$ $0.855$ $0.454$ $16.40$ $0.432$ $12.32$ $0.582$
SegQuant-G $26.76$ $0.843$ $0.438$ $16.45$ $0.434$ $12.40$ $0.578$
SDXL-UNet 8/8(int) Q-Diffusion $74.89$ $-1.574$ $0.396$ $16.48$ $0.537$ $13.68$ $0.517$
PTQ4DiT $7.51$ $0.639$ $0.407$ $16.70$ $0.213$ $19.17$ $0.725$
SegQuant-A $5.88$ $0.645$ $0.408$ $16.71$ $0.145$ $21.32$ $0.785$
SegQuant-G $5.72$ $0.652$ $0.408$ $16.71$ $0.138$ $21.63$ $0.793$
8/8(fp) Q-Diffusion $4.94$ $0.843$ $0.428$ $16.46$ $0.104$ $23.32$ $0.832$
SegQuant1 $4.92$ $0.844$ $0.428$ $16.46$ $0.104$ $23.31$ $0.832$
SegQuant2 $4.63$ $0.839$ $0.427$ $16.46$ $0.093$ $23.90$ $0.842$
FLUX-DiT 8/8(int) Q-Diffusion $8.37$ $0.883$ $0.446$ $16.52$ $0.302$ $15.23$ $0.624$
PTQ4DiT $10.47$ $0.716$ $0.460$ $16.45$ $0.328$ $14.73$ $0.620$
SegQuant-A $5.13$ $0.907$ $0.445$ $16.52$ $0.155$ $19.40$ $0.770$
SegQuant-G $4.94$ $0.900$ $0.444$ $16.54$ $0.143$ $19.94$ $0.784$
4/8(int) PTQ4DiT $34.92$ $0.096$ $0.417$ $16.65$ $0.561$ $11.63$ $0.537$
SVDQuant $7.11$ $0.880$ $0.436$ $16.57$ $0.242$ $16.85$ $0.693$
SegQuant-G $6.86$ $0.882$ $0.439$ $16.56$ $0.232$ $17.03$ $0.700$
@@
SD3.5-DiT 8/8(int) Q-Diffusion $0.444$ $15.31$ 4/8(int) PTQ4DiT 0.428 15.86
PTQD $0.441$ $15.91$
PTQ4DiT $0.461$ $15.91$ SVDQuant 0.452 15.85
TAC $0.460$ $15.90$
SegQuant-A $0.468$ $15.86$
SegQuant-G $0.466$ $15.88$ SegQuant-G 0.452 15.89
SDXL-UNet 8/8(int) Q-Diffusion $0.400$ $15.61$ 8/8(fp) Q-Diffusion 0.417 15.71
PTQ4DiT $0.430$ $15.76$
SegQuant-A $0.433$ $15.77$ SegQuant-A $0.417$ $15.71$
SegQuant-G $0.433$ $15.78$ SegQuant-G $0.418$ $15.71$
FLUX-DiT 8/8(int) Q-Diffusion $0.444$ $15.98$ 4/8(int) PTQ4DiT 0.416 15.98
PTQ4DiT $0.461$ $15.88$
SegQuant-A $0.440$ $15.94$ SVDQuant $0.435$ $15.99$
SegQuant-G $0.440$ $15.93$ SegQuant-G $0.437$ $16.01$
@@
SD3.5-DiT 8/8(int) Q-Diffusion $161.22$ $-1.973$ $0.460$ $17.92$ $0.691$ $7.80$ $0.254$
PTQD $53.38$ $-0.454$ $0.416$ $18.08$ $0.582$ $10.00$ $0.296$
PTQ4DiT $15.71$ $0.485$ $0.430$ $18.08$ $0.445$ $12.33$ $0.492$
TAC $17.22$ $0.430$ $0.432$ $18.10$ $0.461$ $12.11$ $0.478$
SegQuant-A $11.99$ $0.639$ $0.440$ $18.07$ $0.407$ $12.74$ $0.521$
SegQuant-G $11.42$ $0.639$ $0.447$ $18.06$ $0.412$ $12.40$ $0.516$
4/8(int) PTQ4DiT $67.23$ $-0.314$ $0.403$ $18.21$ $0.566$ $10.94$ $0.386$
SVDQuant $20.03$ $0.518$ $0.430$ $18.14$ $0.446$ $12.34$ $0.495$
SegQuant-G $17.16$ $0.576$ $0.435$ $18.14$ $0.438$ $12.25$ $0.497$
SDXL-UNet 8/8(int) Q-Diffusion $64.87$ $-1.466$ $0.402$ $17.65$ $0.552$ $13.64$ $0.448$
PTQ4DiT $7.15$ $0.472$ $0.415$ $17.94$ $0.197$ $19.63$ $0.686$
SegQuant-A $5.67$ $0.487$ $0.416$ $17.94$ $0.132$ $21.69$ $0.752$
SegQuant-G $5.47$ $0.504$ $0.416$ $17.94$ $0.123$ $22.11$ $0.763$
8/8(fp) Q-Diffusion $5.06$ $0.474$ $0.406$ $17.89$ $0.102$ $23.28$ $0.780$
SegQuant-A $4.96$ $0.475$ $0.406$ $17.89$ $0.101$ $23.34$ $0.782$
SegQuant-G $4.77$ $0.474$ $0.405$ $17.89$ $0.092$ $23.82$ $0.793$
FLUX-DiT 8/8(int) Q-Diffusion $8.96$ $0.513$ $0.449$ $18.01$ $0.298$ $14.40$ $0.563$
PTQ4DiT $13.04$ $0.438$ $0.472$ $17.97$ $0.552$ $9.22$ $0.321$
SegQuant-A $5.72$ $0.592$ $0.452$ $18.01$ $0.149$ $18.29$ $0.715$
SegQuant-G $5.32$ $0.590$ $0.451$ $18.02$ $0.131$ $18.97$ $0.737$
4/8(int) PTQ4DiT $60.24$ $-0.393$ $0.418$ $17.88$ $0.592$ $10.98$ $0.465$
SVDQuant $8.12$ $0.575$ $0.442$ $18.03$ $0.224$ $16.19$ $0.635$
SegQuant-G $8.56$ $0.566$ $0.446$ $18.03$ $0.219$ $16.31$ $0.639$

### v3

@@ Introduction
To quantify this asymmetry, we analyzed activation outputs channel-wise across several representative layers from SD3.5-ControlNet using the COCO dataset (averaged over 30 timesteps). As shown in Table 1, a large proportion of channels consistently exhibit negative values. For instance, over 60%-70% of channels in certain AdaNorm and FFN layers remain predominantly negative during inference. However, conventional PTQ methods typically apply a single global scale (either symmetric or asymmetric) across the entire activation range. Due to the large spread of positive values, compression of the narrower negative range ([0.3,0][-0.3,0][ - 0.3 , 0 ]) is often too aggressive. Even with asymmetric quantization, which shifts the zero-point, the uniform distribution of bins leads to poor resolution for negative values. Figure 7 shows that modifying only the negative activations causes subtle but perceptible detail loss, particularly in texture and background consistency. This indicates that negative activations are not residual noise but semantically meaningful components. Although overall quality remains largely intact, this underscores preserving small negative values for high-fidelity generation.
@@
SD3.5-DiT 8/8(int) Q-Diffusion 169.59169.59169.59 2.072-2.072- 2.072 0.6900.6900.690 7.927.927.92 0.2950.2950.295
PTQD 26.5326.5326.53 0.3090.3090.309 0.5200.5200.520 10.2010.2010.20 0.4170.4170.417
PTQ4DiT 16.4616.4616.46 0.7520.7520.752 0.4260.4260.426 12.1812.1812.18 0.5320.5320.532
TAC 17.7817.7817.78 0.7020.7020.702 0.4400.4400.440 11.9911.9911.99 0.5200.5200.520
SegQuant-A 13.9013.9013.90 0.9240.9240.924 0.3840.3840.384 12.7812.7812.78 0.5630.5630.563
SegQuant-G 12.3712.3712.37 0.8590.8590.859 0.3830.3830.383 12.8312.8312.83 0.5640.5640.564
4/8(int) PTQ4DiT 62.9862.9862.98 0.190-0.190- 0.190 0.5770.5770.577 10.0610.0610.06 0.4290.4290.429
SVDQuant 20.5820.5820.58 0.7250.7250.725 0.4560.4560.456 11.7611.7611.76 0.5230.5230.523
SegQuant-G 20.2220.2220.22 0.7620.7620.762 0.4530.4530.453 11.6911.6911.69 0.5210.5210.521
SDXL-UNet 8/8(int) Q-Diffusion 72.8172.8172.81 1.618-1.618- 1.618 0.5670.5670.567 11.9711.9711.97 0.3600.3600.360
PTQ4DiT 7.707.707.70 0.7360.7360.736 0.1910.1910.191 19.6619.6619.66 0.6910.6910.691
SegQuant-A 6.396.396.39 0.7750.7750.775 0.1410.1410.141 21.3321.3321.33 0.7420.7420.742
SegQuant-G 6.196.196.19 0.7640.7640.764 0.1340.1340.134 21.6021.6021.60 0.7500.7500.750
8/8(fp) Q-Diffusion 5.145.145.14 0.8970.8970.897 0.0930.0930.093 24.3124.3124.31 0.8270.8270.827
SegQuant-A 5.125.125.12 0.9010.9010.901 0.0930.0930.093 24.2824.2824.28 0.8270.8270.827
SegQuant-G 4.834.834.83 0.9030.9030.903 0.0820.0820.082 24.8424.8424.84 0.8380.8380.838
FLUX-DiT 8/8(int) Q-Diffusion 9.419.419.41 0.7320.7320.732 0.2990.2990.299 15.8715.8715.87 0.6330.6330.633
PTQ4DiT 11.9111.9111.91 0.6300.6300.630 0.3250.3250.325 15.3615.3615.36 0.6110.6110.611
SegQuant-A 5.895.895.89 0.8350.8350.835 0.1500.1500.150 19.8419.8419.84 0.7700.7700.770
SegQuant-G 5.565.565.56 0.8220.8220.822 0.1380.1380.138 20.3220.3220.32 0.7820.7820.782
4/8(int) PTQ4DiT 30.0930.0930.09 0.039-0.039- 0.039 0.5400.5400.540 12.0912.0912.09 0.5400.5400.540
SVDQuant 7.947.947.94 0.7830.7830.783 0.2320.2320.232 17.2917.2917.29 0.6970.6970.697
SegQuant-G 7.787.787.78 0.7890.7890.789 0.2250.2250.225 17.4817.4817.48 0.7020.7020.702
@@ Appendix B More Experimental Details and Results
SD3.5-DiT 8/8(int) Q-Diffusion 218.04218.04218.04 2.197-2.197- 2.197 0.3950.3950.395 16.0216.0216.02 0.7040.7040.704 8.218.218.21 0.3880.3880.388
PTQD 15.6215.6215.62 0.8220.8220.822 0.4340.4340.434 16.4716.4716.47 0.4540.4540.454 10.7910.7910.79 0.5230.5230.523
PTQ4DiT 14.2114.2114.21 0.9120.9120.912 0.4580.4580.458 16.3716.3716.37 0.4030.4030.403 12.6912.6912.69 0.5930.5930.593
TAC 16.8916.8916.89 0.8630.8630.863 0.4580.4580.458 16.3116.3116.31 0.4170.4170.417 12.5312.5312.53 0.5830.5830.583
SegQuant-A 10.9010.9010.90 1.0201.0201.020 0.4670.4670.467 16.3516.3516.35 0.3620.3620.362 13.2713.2713.27 0.6180.6180.618
SegQuant-G 11.0611.0611.06 0.9910.9910.991 0.4570.4570.457 16.3616.3616.36 0.3760.3760.376 13.1413.1413.14 0.6010.6010.601
4/8(int) PTQ4DiT 92.3692.3692.36 0.0850.0850.085 0.4110.4110.411 16.5116.5116.51 0.5700.5700.570 10.8310.8310.83 0.5040.5040.504
SVDQuant 25.4825.4825.48 0.8550.8550.855 0.4540.4540.454 16.4016.4016.40 0.4320.4320.432 12.3212.3212.32 0.5820.5820.582
SegQuant-G 26.7626.7626.76 0.8430.8430.843 0.4380.4380.438 16.4516.4516.45 0.4340.4340.434 12.4012.4012.40 0.5780.5780.578
SDXL-UNet 8/8(int) Q-Diffusion 74.8974.8974.89 1.574-1.574- 1.574 0.3960.3960.396 16.4816.4816.48 0.5370.5370.537 13.6813.6813.68 0.5170.5170.517
PTQ4DiT 7.517.517.51 0.6390.6390.639 0.4070.4070.407 16.7016.7016.70 0.2130.2130.213 19.1719.1719.17 0.7250.7250.725
SegQuant-A 5.885.885.88 0.6450.6450.645 0.4080.4080.408 16.7116.7116.71 0.1450.1450.145 21.3221.3221.32 0.7850.7850.785
SegQuant-G 5.725.725.72 0.6520.6520.652 0.4080.4080.408 16.7116.7116.71 0.1380.1380.138 21.6321.6321.63 0.7930.7930.793
8/8(fp) Q-Diffusion 4.944.944.94 0.8430.8430.843 0.4280.4280.428 16.4616.4616.46 0.1040.1040.104 23.3223.3223.32 0.8320.8320.832
SegQuant1 4.924.924.92 0.8440.8440.844 0.4280.4280.428 16.4616.4616.46 0.1040.1040.104 23.3123.3123.31 0.8320.8320.832
SegQuant2 4.634.634.63 0.8390.8390.839 0.4270.4270.427 16.4616.4616.46 0.0930.0930.093 23.9023.9023.90 0.8420.8420.842
FLUX-DiT 8/8(int) Q-Diffusion 8.378.378.37 0.8830.8830.883 0.4460.4460.446 16.5216.5216.52 0.3020.3020.302 15.2315.2315.23 0.6240.6240.624
PTQ4DiT 10.4710.4710.47 0.7160.7160.716 0.4600.4600.460 16.4516.4516.45 0.3280.3280.328 14.7314.7314.73 0.6200.6200.620
SegQuant-A 5.135.135.13 0.9070.9070.907 0.4450.4450.445 16.5216.5216.52 0.1550.1550.155 19.4019.4019.40 0.7700.7700.770
SegQuant-G 4.944.944.94 0.9000.9000.900 0.4440.4440.444 16.5416.5416.54 0.1430.1430.143 19.9419.9419.94 0.7840.7840.784
4/8(int) PTQ4DiT 34.9234.9234.92 0.0960.0960.096 0.4170.4170.417 16.6516.6516.65 0.5610.5610.561 11.6311.6311.63 0.5370.5370.537
SVDQuant 7.117.117.11 0.8800.8800.880 0.4360.4360.436 16.5716.5716.57 0.2420.2420.242 16.8516.8516.85 0.6930.6930.693
SegQuant-G 6.866.866.86 0.8820.8820.882 0.4390.4390.439 16.5616.5616.56 0.2320.2320.232 17.0317.0317.03 0.7000.7000.700
@@
SD3.5-DiT 8/8(int) Q-Diffusion 0.4440.4440.444 15.3115.3115.31 4/8(int) PTQ4DiT 0.428 15.86
PTQD 0.4410.4410.441 15.9115.9115.91
PTQ4DiT 0.4610.4610.461 15.9115.9115.91 SVDQuant 0.452 15.85
TAC 0.4600.4600.460 15.9015.9015.90
SegQuant-A 0.4680.4680.468 15.8615.8615.86
SegQuant-G 0.4660.4660.466 15.8815.8815.88 SegQuant-G 0.452 15.89
SDXL-UNet 8/8(int) Q-Diffusion 0.4000.4000.400 15.6115.6115.61 8/8(fp) Q-Diffusion 0.417 15.71
PTQ4DiT 0.4300.4300.430 15.7615.7615.76
SegQuant-A 0.4330.4330.433 15.7715.7715.77 SegQuant-A 0.4170.4170.417 15.7115.7115.71
SegQuant-G 0.4330.4330.433 15.7815.7815.78 SegQuant-G 0.4180.4180.418 15.7115.7115.71
FLUX-DiT 8/8(int) Q-Diffusion 0.4440.4440.444 15.9815.9815.98 4/8(int) PTQ4DiT 0.416 15.98
PTQ4DiT 0.4610.4610.461 15.8815.8815.88
SegQuant-A 0.4400.4400.440 15.9415.9415.94 SVDQuant 0.4350.4350.435 15.9915.9915.99
SegQuant-G 0.4400.4400.440 15.9315.9315.93 SegQuant-G 0.4370.4370.437 16.0116.0116.01
@@
SD3.5-DiT 8/8(int) Q-Diffusion 161.22161.22161.22 1.973-1.973- 1.973 0.4600.4600.460 17.9217.9217.92 0.6910.6910.691 7.807.807.80 0.2540.2540.254
PTQD 53.3853.3853.38 0.454-0.454- 0.454 0.4160.4160.416 18.0818.0818.08 0.5820.5820.582 10.0010.0010.00 0.2960.2960.296
PTQ4DiT 15.7115.7115.71 0.4850.4850.485 0.4300.4300.430 18.0818.0818.08 0.4450.4450.445 12.3312.3312.33 0.4920.4920.492
TAC 17.2217.2217.22 0.4300.4300.430 0.4320.4320.432 18.1018.1018.10 0.4610.4610.461 12.1112.1112.11 0.4780.4780.478
SegQuant-A 11.9911.9911.99 0.6390.6390.639 0.4400.4400.440 18.0718.0718.07 0.4070.4070.407 12.7412.7412.74 0.5210.5210.521
SegQuant-G 11.4211.4211.42 0.6390.6390.639 0.4470.4470.447 18.0618.0618.06 0.4120.4120.412 12.4012.4012.40 0.5160.5160.516
4/8(int) PTQ4DiT 67.2367.2367.23 0.314-0.314- 0.314 0.4030.4030.403 18.2118.2118.21 0.5660.5660.566 10.9410.9410.94 0.3860.3860.386
SVDQuant 20.0320.0320.03 0.5180.5180.518 0.4300.4300.430 18.1418.1418.14 0.4460.4460.446 12.3412.3412.34 0.4950.4950.495
SegQuant-G 17.1617.1617.16 0.5760.5760.576 0.4350.4350.435 18.1418.1418.14 0.4380.4380.438 12.2512.2512.25 0.4970.4970.497
SDXL-UNet 8/8(int) Q-Diffusion 64.8764.8764.87 1.466-1.466- 1.466 0.4020.4020.402 17.6517.6517.65 0.5520.5520.552 13.6413.6413.64 0.4480.4480.448
PTQ4DiT 7.157.157.15 0.4720.4720.472 0.4150.4150.415 17.9417.9417.94 0.1970.1970.197 19.6319.6319.63 0.6860.6860.686
SegQuant-A 5.675.675.67 0.4870.4870.487 0.4160.4160.416 17.9417.9417.94 0.1320.1320.132 21.6921.6921.69 0.7520.7520.752
SegQuant-G 5.475.475.47 0.5040.5040.504 0.4160.4160.416 17.9417.9417.94 0.1230.1230.123 22.1122.1122.11 0.7630.7630.763
8/8(fp) Q-Diffusion 5.065.065.06 0.4740.4740.474 0.4060.4060.406 17.8917.8917.89 0.1020.1020.102 23.2823.2823.28 0.7800.7800.780
SegQuant-A 4.964.964.96 0.4750.4750.475 0.4060.4060.406 17.8917.8917.89 0.1010.1010.101 23.3423.3423.34 0.7820.7820.782
SegQuant-G 4.774.774.77 0.4740.4740.474 0.4050.4050.405 17.8917.8917.89 0.0920.0920.092 23.8223.8223.82 0.7930.7930.793
FLUX-DiT 8/8(int) Q-Diffusion 8.968.968.96 0.5130.5130.513 0.4490.4490.449 18.0118.0118.01 0.2980.2980.298 14.4014.4014.40 0.5630.5630.563
PTQ4DiT 13.0413.0413.04 0.4380.4380.438 0.4720.4720.472 17.9717.9717.97 0.5520.5520.552 9.229.229.22 0.3210.3210.321
SegQuant-A 5.725.725.72 0.5920.5920.592 0.4520.4520.452 18.0118.0118.01 0.1490.1490.149 18.2918.2918.29 0.7150.7150.715
SegQuant-G 5.325.325.32 0.5900.5900.590 0.4510.4510.451 18.0218.0218.02 0.1310.1310.131 18.9718.9718.97 0.7370.7370.737
4/8(int) PTQ4DiT 60.2460.2460.24 0.393-0.393- 0.393 0.4180.4180.418 17.8817.8817.88 0.5920.5920.592 10.9810.9810.98 0.4650.4650.465
SVDQuant 8.128.128.12 0.5750.5750.575 0.4420.4420.442 18.0318.0318.03 0.2240.2240.224 16.1916.1916.19 0.6350.6350.635
SegQuant-G 8.568.568.56 0.5660.5660.566 0.4460.4460.446 18.0318.0318.03 0.2190.2190.219 16.3116.3116.31 0.6390.6390.639